# TRecViT: A Recurrent Video Transformer

### v1

@@ 5.1 Supervised video classification
Figure 4: TRecViT compared to baselines on supervised video classification on SSv2 dataset, trained from scratch. The plot shows the evolution of the evaluation accuracy as training progresses.
Figure 5: Qualitative results obtained by TRecViT for point tracking on DAVIS dataset compared to VideoMAE. The leftmost image indicates the point to track in the original frame, and the images towards the right show zoom-ins on subsequent frames. Green plus (+) marker indicates the ground truth, yellow circle indicates TRecViT’s predictions and red circles indicate VideoMAE’s predictions.
Model Patch size Top-1 acc (%) # params
ViViT-B (2, 16, 16) 59.1 90M
ViViT-L (2, 16, 16) 65.9 320M
TRecViT (1, 16, 16) 66.8 109M
Table 1: Performance of TRecViT compared to ViViT-B and ViViT-L baselines on SSv2 dataset with all models initialised from Imagenet pre-training. For ViViT-L, we use the result reported by its authors, for ViViT-B we obtained the results internally as they were not reported in the original paper for this dataset.
Model Patch size Top-1 acc (%) # params
ViViT-B (2, 16, 16) 78.1 90M
ViViT-L (2, 16, 16) 78.7 320M
TRecViT (1, 16, 16) 78.4 109M
Table 2: Performance of TRecViT compared to ViViT-B and ViViT-L baselines on Kinetics400 dataset, with all models initialised from Imagenet pre-training. For ViViT-B and ViViT-L, we include the result we obtained internally by re-training the model on the current Kinetics400 dataset version; see footnote. In the original paper, the authors reported 80.3% on Kinetics400 for ViViT-L.
@@ 5.2 Self-supervised masked autoencoding
Video classification: We report video classification accuracy as downstream task using attention readout heads on SSv2 and Kinetics400. We compare the performance against VideoMAE-L [40] in Table 3. Our model obtains slightly better performance on both datasets compared to this strong baseline, despite having almost 3$\times$ less parameters.
Point tracking: To demonstrate that our model can handle dense(r) tasks as well, we evaluate the same frozen MAE representations for the point tracking task. We use the recurrent architecture in MooG [41] as a readout due to its simplicity. MooG uses light cross-attention layers to process the embeddings of each frame in order, and the readout state is carried over through time. We finetune the MooG readout head using MOVi-E dataset [25] as done in popular point tracking works [11]. We evaluate these fine-tuned representations on two datasets: Perception Test [36] and DAVIS dataset [35] with point tracks extracted in [10]. We report average Jaccard metric [10] for TRecViT compared with MooG and VideoMAE; see Table 4. TRecViT obtains better performance on both datasets compared to baselines, which reinforces the observation that our proposed model has strong motion modelling capabilities. We include qualitative results for this task in Figure 5. We can observe that the results are visibly better compared to VideoMAE. More visualisations are included in the supplementary material.
Model Dataset Top-1 acc (%) # params
VideoMAE Kinetics400 45.8 330M
TRecViT Kinetics400 46.0 128M
VideoMAE SSv2 53.7 330M
TRecViT SSv2 53.9 128M
Table 3: Performance of TRecViT compared to VideoMAE on video classification using frozen MAE representations, pre-trained on Kinetics400.
Model Dataset # frames AJ # params
MooG DAVIS 8 0.687 35M
VideoMAE DAVIS 8 0.703 330M
TRecViT DAVIS 8 0.706 128M
MooG Perception Test 16 0.760 46.5M
VideoMAE Perception Test 16 0.761 330M
TRecViT Perception Test 16 0.783 128M
Table 4: Performance of TRecViT compared to baselines on point tracking task on DAVIS and Perception Test datasets. All models use frozen representations evaluated using the readout head from MooG.
Figure 6: Qualitative results obtained by TRecViT on the dense memorisation task compared to ViViT-L. Both models are trained using Imagenet pre-trained weights, on video sequences of $T=64$ frames and they reconstruct the $(T-48)^{\text{th}}$ frame.
(a) PSNR comparison
(b) Steps-per-second comparison
Figure 7: Long video memorisation task. At time $T$, the model has to reconstruct the $(T-k)^{\text{th}}$ frame seen in the past. The plots show PSNR and throughput (steps-per-second) for increasing time offset $k$. For both models, the data points with 00 value on the $y$-axis correspond to OOM.
@@ 5.3 Long video memorisation task
Transformer models for language are known to be excellent at retrieving information from context, as they cache the keys and values for the entire history. On the other hand, LRUs / SSMs and RNNs in general struggle with such needle-in-the-haystack style tasks as they need to perform the retrieval based on the compressed history kept in their recurrent state [24, 9]. We are interested in studying this aspect in the video domain as well. We set up a simple reconstruction task where the model has to remember the frame seen at a given time-step in the past. For our analysis, we run multiple experiments where the model is tasked to reconstruct the $(T-k)^{\text{th}}$ frame from the past, with increasing value for $k\in\{16,48,80,112,144,164\}$ frames. We employ Walking Tours dataset [43], which contains hour-long videos, and the scenery changes constantly, hence we are guaranteed that the video frames seen most recently will be very different compared to the frames seen earlier on. We scale the videos to $224\times 224$ pixels. Again, we adopt ViViT-L as baseline, and we train both models using Imagenet pretrained weights. For ViViT-L, we keep all the outputs from all $T$ time steps and apply temporal pooling and a $1\times 1$ convolution to get the expected shape for the reconstructed frame. For TRecViT, we simply keep the output of the last layer at time step $T$ and reshape it to the expected shape. We show quantitative and qualitative results respectively in Figures 7 and 6. We can observe that there is a performance–efficiency trade-off at play for TRecViT: its performance is slightly below ViViT’s for shorter memory spans (16, 48, 80), but its efficiency (steps-per-second) is significantly higher. However, beyond 80 frames, ViViT-L goes out of memory, whilst TRecViT continues to give decent results up to 144 frames, going out of memory towards 164 frames. Figure 6 shows qualitative results compared to the baseline for the case where the models have to remember the frame seen at $T-48$ in the past. We can observe that the quality of ViViT-L’s reconstruction is good. For TRecViT, whilst the overall structure (encoded in lower frequencies) is correct, it struggles to remember the high-frequency content of the image. This is to be expected due to the compression happening in the recurrent state of the model. However, given how different the last seen frame is from the target frame, we consider this to be a very promising result that warrants further investigation into the memorisation capabilities of our model, which we leave as future work.
5.4 Generalisation to longer sequences
Using the same task as above, we analyse the generalisation capabilities to sequences longer than those used during training. Specifically, we train the models with sequences of length $T=64$ frames to reconstruct the $T-48$ frame, and evaluate them on longer sequences $T=96$ to reconstruct the same frame. The TRecViT model can run on longer sequences without any modification. For the ViViT model, we need to adapt the positional encoding to accommodate longer sequences. We use interpolation to nearest neighbour to obtain the desired length; cubic interpolation led to worse results. The performance of TRecViT degrades slightly, with PSNR going down from 29.3 (when evaluated on the same sequence length as in training $T=64$) to 26.4 when evaluated with $T=96$ frame sequences. ViViT’s PSNR, however, drops significantly, from 32.3 when evaluated on the same sequence length, to 15.1 when evaluated on longer sequences. We include qualitative examples in Figure 8 where we can observe that ViViT’s output contains stronger artefacts compared to TRecViT.
Figure 8: Generalisation to longer sequences. Both models are trained using Imagenet pre-trained weights, on video sequences of $T=64$ frames to reconstruct the $(T-48)^{\text{th}}$ frame; during evaluation, the models receive sequences of $T=96$ frames.
@@ 6 Conclusion
We propose a novel video architecture TRecViT that alternates gated linear recurrent units (LRUs) modelling the temporal dynamics in the video with ViT blocks modelling the spatial and channel dimensions. The proposed model outperforms or obtains competitive performance compared to strong baselines (ViViT-L, VideoMAE) on supervised and self-supervised tasks, while having a much smaller number of parameters and significantly reduced memory footprint and FLOPs count. In terms of limitations, our study focuses on doing a first investigation into using LRUs for the video domain and we obtain favourable results on multiple datasets and tasks compared to strong baselines. However, more experimentation and model scaling are required to obtain SOTA results on all these tasks. Given that the training dynamics for gated LRUs are stable and controllable by design, plus the reliance on (pre-trained) ViT blocks give a strong indication that achieving SOTA is possible. We leave this investigation for future work, together with further analysis of training dynamics, and integration into various downstream tasks, e.g. video-language tasks or Robotics tasks.
Acknowledgements
We would like to thank Caglar Gulcehre, Daniel Zoran, Dima Damen, and Andrew Zisserman for their insightful feedback throughout this project.
Supplementary Material
We include here all the hyperparameters used in the experiments presented in the main paper, together with more qualitative visualisations of results for the point tracking task (section 5.2) and the long video memorisation task (section 5.3). Videos showing point tracks are also attached.
7 Training hyperparameters
7.1 Supervised video classification
@@
Table 5: Hyperparameter values used in the supervised classification experiments. These are mainly the hyperparameters used in previous works, e.g. ViViT [1]. For both datasets, we use cosine decay for the learning rate schedule with linear warmup.
7.2 Self-supervised masked autoencoding and fine-tuning
@@
Table 6: Hyperparameter values used in the self-supervised masked auto-encoding experiment on Kinetics400. We use AdamW optimizer. We apply patch-wise normalisation of the inputs as done in VideoMAE [40]
@@
8 Point tracking qualitative results
In Figure 9, we include more visualisations for the point tracking task using frozen MAE representations pre-trained on Kinetics400, using TRecViT as backbone.
@@ 8 Point tracking qualitative results
9 Long video memorisation task
Figure 10 shows qualitative results for the memorisation task. For easier visual comparison, we increase the distance $k$ to the frame to reconstruct while also increasing the video length $T$, so the frame to reconstruct is always the same. For ViViT-L (3rd row), the quality of the reconstruction is very good and does not degrade as $k$ increases. However, the model goes out-of-memory for $T>96$. For TRecViT, the high frequencies are less well reconstructed as $k$ increases, but overall the model is able to perform the task reasonably well even at $T=160,k=144$, i.e. it is able to learn with sequences of up to 5.3s long at 30FPS, and remember a frame seen about 4.8s before.

### v2

@@ 5.2 SOTA comparison on supervised video classification
Figure 4: TRecViT compared to baselines on supervised video classification on SSv2 dataset, trained from scratch. The plot shows the evolution of the evaluation accuracy as training progresses (higher is better).
Model Pre- Top-1 Param FLOPs Mem
Train (%) (M) (T) (G)
non-causal
I3D K400 51.3 25.0 N/A N/A
SlowFast R50 K400 61.9 34.1 0.19 3.35
SlowFast R101 K400 63.1 53.3 0.32 4.20
MSNet IN-21K 64.7 54.6 0.07 6.54
TimeSformer-L IN-21K 62.4 121.4 5.1 >>24
VidTr-L - 63.0 N/A 10.5 N/A
ViViT-L32 - 65.9 310.8 7.75 6.11
Mformer-B IN-21K 66.5 114.0 1.10 7.3
MViT-B32 K400 67.1 36.6 0.51 10.7
MViT-B64 K400 67.7 36.6 1.36 >>24
MViT-B-2432 K600 68.7 36.6 1.36 >>24
VideoMamba IN-21K 68.4 74.0 0.33 N/A
VideoMambaPro IN-21K 69.4 72.0 2.2 N/A
causal
TSM* K400 63.3 42.9 0.19 5.98
cViViT-L32 IN-21K 64.4 310.8 N/A N/A
RViT-L32 K400 66.1 72.0 1.34 2.12
RVIT-XL64 K400 67.9 107.7 3.99 2.33
TRecViT32 IN-21K 66.8 111.3 1.44 1.79
TRecViT64 IN-21K 68.2 111.3 2.89 3.16**
Table 3: Performance comparison on SSv2 dataset, considering causal and non-causal baselines. cViViT-L is a causal ViViT model that we trained using causal attention masking for the self-attention block. We do not include FLOPs and Mem for this model as they will heavily depend on the implementation of the causal masking. Where present, the subscript indicates the number of frames used at inference time. Metric: Top-1 accuracy (higher is better). *Result reported in (Yang et al., 2022). **For TRecViT, the memory at inference is independent of the number of frames, here we run causally on the entire sequence in parallel, hence the memory increase.
Model Pre-Train Top-1 (%) Param (M)
non-causal
I3D IN-1K 72.1 25.0
TimeSformer IN-21K 78.0 121.4
Mformer-B IN-21K 79.7 114.0
ViViT-L IN-21K 80.3 320.0
ViViT-B IN-21K 78.1 91.2
ViViT-L IN-21K 78.7 310.9
causal
RViT-XL IN-21K 81.5 107.7
cViViT-L IN-21K 76.3 310.9
TRecViT IN-21K 76.5 111.2
Table 4: Performance of TRecViT compared to convolutional and transformer-based baselines on Kinetics400 dataset. For fair comparison, we trained ViViT-B and ViViT-L models on the current Kinetics400 dataset version; see footnote 2. The numbers in gray correspond to results obtained on the original larger Kinetics400 dataset, reported by their authors. Metric: Top-1 accuracy (higher is better).
@@ 5.3 Self-supervised masked autoencoding
[…]
Model Dataset Top-1 (%) Param (M)
VideoMAE Kinetics400 45.8 330
TRecViT Kinetics400 46.0 128
VideoMAE SSv2 53.7 330
TRecViT SSv2 53.9 128
Table 5: Performance of TRecViT compared to VideoMAE on video classification using frozen MAE representations, pre-trained on Kinetics400. Metric: Top-1 accuracy (higher is better).
Point tracking: To demonstrate that our model can handle dense(r) tasks as well, we evaluate the same frozen MAE representations for the point tracking task. We use the recurrent architecture in MooG (van Steenkiste et al., 2024) as a readout head due to its simplicity. MooG uses light cross-attention layers to process the embeddings of each frame in order, and the readout state is carried over time. We finetune the readout head using MOVi-E dataset (Greff et al., 2022) as done in popular point tracking works (Doersch et al., 2023). We evaluate these fine-tuned representations on two datasets: Perception Test (Pătrăucean et al., 2023) and DAVIS dataset (Pont-Tuset et al., 2017) with point tracks extracted in (Doersch et al., 2022). We report average Jaccard metric, i.e. the percentage of predicted points that fall within a certain threshold from the ground truth points (Doersch et al., 2022), for TRecViT compared with MooG and VideoMAE; see Table 7. TRecViT obtains better performance on both datasets compared to baselines, which reinforces the observation that our proposed model has strong motion modelling capabilities. We include qualitative results for this task in the appendix.
@@ 5.4 Long video memorisation task
Transformer models for language are known to be excellent at retrieving information from context, as they cache the keys and values for the entire history. On the other hand, LRUs / SSMs and RNNs in general struggle with such needle-in-the-haystack style tasks as they need to perform the retrieval based on the compressed history kept in their recurrent state (Jelassi et al., 2024; De et al., 2024).
We are interested in studying this aspect in the video domain as well. We set up a simple reconstruction task where the model has to remember the frame seen at a given time-step in the past. For example, we train the models with sequences of length $T=96$ frames to reconstruct the 16th frame. Using the same task, we also analyse the generalisation capabilities to sequences longer than those seen during training, i.e. the models are trained with sequences of length $T=64$ frames to reconstruct the 16th frame, and then are evaluated on video sequences with $T=96$ frames to reconstruct the same 16th frame.
Model Num frames train Num frames test PSNR
ViViT-L 96 96 32.2
TRecViT 96 96 29.1
ViViT-L 64 96 15.1
TRecViT 64 96 26.4
Table 6: Comparison between TRecViT and ViViT-L on the long video memorisation task, when the models are evaluated on the same number of frames as seen in training (top) or on longer sequences (bottom). Metric: Peak Signal-To-Noise Ratio (PSNR); higher is better.
We employ Walking Tours dataset (Venkataramanan et al., 2024), which contains hour-long videos, and the scenery changes constantly, hence we are guaranteed that the video frames seen most recently will be very different compared to the frames seen earlier on. We scale the videos to $224\times 224$ pixels. We adopt ViViT-L as baseline, and we train both models using Imagenet pretrained weights. For ViViT-L, we keep all the outputs from all $T$ time steps and apply temporal pooling and a $1\times 1$ convolution to get the expected shape for the reconstructed frame. For TRecViT, we simply keep the output of the last layer at time step $T$ and reshape it to the expected shape. When evaluating on longer sequences, the ViViT model needs to adapt the positional encoding – we use interpolation to nearest neighbour to obtain the desired length; cubic interpolation led to worse results. TRecViT can run on any sequence length without modification.
Model Dataset AJ Param (M)
MooG DAVIS 0.687 35
VideoMAE DAVIS 0.703 330
TRecViT DAVIS 0.706 128
MooG Perception Test 0.760 46.5M
VideoMAE Perception Test 0.761 330M
TRecViT Perception Test 0.783 128M
Table 7: Performance of TRecViT compared to baselines on point tracking task on DAVIS and Perception Test datasets. All models use frozen representations evaluated using the readout head from MooG. In line with MooG, we use 8 frames for the DAVIS dataset and 16 frames for the Perception Test. Metric: average Jaccard (AJ); higher is better.
Table 6 includes quantitative results measured in terms of Peak Signal-to-Noise Ratio (PSNR), which is defined based on the mean squared error (MSE) between the pixel values of the original video and the reconstructed video; more quantitative and qualitative results are included in the Appendix. When evaluated on the same sequence length, ViViT-L is better than TRecViT as it manages to better preserve the higher frequencies in the visual signal. However, when evaluated on longer sequences, ViViT’s PSNR drops significantly, showing strong artefacts, whilst TRecViT’s output quality remains quite satisfactory (see qualitative results in the appendix, Figures 8 and 9). We consider this to be a very promising result that warrants further investigation into the memorisation capabilities of our model – we leave this as future work.
@@ 6 Conclusion
We propose a novel causal video architecture TRecViT that alternates gated linear recurrent units (LRUs) modelling the temporal dynamics in the video with ViT blocks modelling the spatial and channel dimensions. Notably, this is the first causal architecture in the SSM video models family. The proposed model outperforms or obtains competitive performance compared to strong causal and non-causal baselines on supervised and self-supervised tasks, while having a much smaller number of parameters and significantly reduced memory footprint and FLOPs count compared to vanilla Transformer video models. Our study indicates that temporal linear recurrence combined with spatial self-attention is a natural and effective parameterisation for video modelling, given the sequential nature of videos as defined by the arrow-of-time.
Limitations and Future Work: While this study presents the first exploration into leveraging LRUs for the video domain, some key aspects necessitate further investigation. First, we plan to conduct a deeper investigation into the memorisation limitations of our architecture, particularly in needle-in-haystack problems. Second, we aim to understand our model’s capabilities in a multi-step generative context, specifically by integrating it within video diffusion models. Finally, given LRUs’ inspiration from continuous time systems, we plan to explore its capabilities of modelling variable frame rate videos across various complex applications, including video-language tasks and Robotics.
Acknowledgments
We would like to thank Çağlar Gülçehre, Jörg Bornschein, Zhitao Gong, Daniel Zoran, and Andrew Zisserman for providing insightful feedback on this work.
Appendix A Model configurations
Table 8 includes the model configurations used in our experiments.
Model Layers Hidden size $D$ ViT MLP size ViT Heads
TRecViT-Small 12 384 1536 6
TRecViT-Base 12 768 3072 12
TRecViT-Large 24 1024 4096 16
Table 8: Model Configurations used in our experiments
Appendix B Training hyperparameters
B.1 Supervised video classification
@@
Table 9: Hyperparameter values used in the supervised classification experiments. These are mainly the hyperparameters used in previous works, e.g. ViViT Arnab et al. (2021). For both datasets, we use cosine decay for the learning rate schedule with linear warmup.
B.2 Self-supervised masked autoencoding and fine-tuning
@@
Table 10: Hyperparameter values used in the self-supervised masked auto-encoding experiment on Kinetics400. We use AdamW optimizer. We apply patch-wise normalisation of the inputs as done in VideoMAE Tong et al. (2022)
@@
Appendix C Efficiency comparison against ViViT variants and other hybrid baselines
Figure 5 complements Figure 3 from the main paper and includes the memory footprint and FLOP counts for other hybrid baselines that use other types of temporal modules (Conv1D, LSTM) instead of the gated LRU used in our proposed architecture. As expected, these hybrid architectures scale linearly in the number of frames, similarly to our model. However, they are outperformed by our model, as mentioned in Table 2 in the main paper.
Figure 5: Left: Memory comparison; Right: FLOPs comparison. Our model has similar efficiency to other factorised architectures, like ViViT factorised self-attention, ViViT factorised encoder, ViT-Conv1D, ViT-LSTM, but it outperforms these baselines in accuracy and generality.
Appendix D Ablations
Patch temporal size Top-1 (%)
1 66.8
2 64.5
4 61.5
8 57.7
Table 13: Performance when using different temporal sizes for the video patches for supervised classification on SSv2, using 32 frames per clip. Our model performs best when the input is fed as spatial patches $t=1$, with accuracy dropping significantly when using $t>1$. We hypothesise that an increased temporal size leads to a less continuous signal fed into the LRUs, affecting its performance.
Window size Top-1 (%)
2 66.4
4 66.8
8 65.7
Table 14: Performance when using different window sizes for the conv 1D kernel in the LRU, for supervised classification on SSv2, using 32 frames per clip. As found in Griffin as well, the best window size is 4.
We include here ablations for different hyperparameters used in our model by running supervised classification experiments on SSv2. We sweep the following hyperparameters: temporal size of the video patches (Table 13), window of the 1D convolution kernel applied in LRU (Table 14), value of the minimal radius when initialising the eigenvalues of the recurrence matrix (Table 15). Finally, we run an experiment using five seeds on SSv2 classification using 32 frames, obtaining $66.6\pm 0.2$; we include the best seed result (seed=0) in our SOTA comparison as done in other works as well (Yang et al., 2022; Arnab et al., 2021), and use this seed for the other experiments.
Min rad eigenvalues Top-1 (%)
0.6 66.8
0.7 66.6
0.8 66.5
0.9 66.2
Table 15: Performance when using different values for the minimal radius when initialising the eigenvalues of the recurrence matrix for supervised classification on SSv2, using 32 frames per clip. Compared to Griffin where $\lambda_{\min}=0.9$ was found to give the best results, for video it is important to lower this value to 0.6, to allow for faster decay of information for some frequencies. We plan to conduct more investigations on this aspect to better understand the connection between $\lambda_{\min}$ and the temporal context of the task being performed.
Appendix E Point tracking qualitative results
In Figure 6, we include more visualisations for the point tracking task using frozen MAE representations pre-trained on Kinetics400, using TRecViT as backbone. Videos showing point tracks are also attached.
@@ Appendix E Point tracking qualitative results
Appendix F Long video memorisation task
We run multiple experiments where the model is tasked to reconstruct the $(T-k)^{\text{th}}$ frame from the past, with increasing value for $k\in\{16,48,80,112,144,164\}$ frames. For easier visual comparison, we increase the distance $k$ to the frame to reconstruct while also increasing the video length $T$, so the frame to reconstruct is always the same.
We show quantitative and qualitative results in Figure 7 and Figure 8, respectively. We can observe that there is a performance–efficiency trade-off at play for TRecViT: its performance is slightly below ViViT’s for shorter memory spans (16, 48, 80), with the high frequencies being less well reconstructed as $k$ increases, but its efficiency (steps-per-second) is significantly higher. However, beyond 80 frames, ViViT-L goes out of memory being unusable, whilst TRecViT continues to give decent results up to $T=160,k=144$, i.e. it is able to learn with sequences of up to 5.3s long at 30FPS, and remember a frame seen about 4.8s before.
Figure 7: Long video memorisation task. Left: PSNR comparison; Right: Step-per-second comparison. At time $T$, the model has to reconstruct the $(T-k)^{\text{th}}$ frame seen in the past. The plots show PSNR and throughput (steps-per-second) for increasing time offset $k$. For both models, the data points with 0 value on the $y$-axis correspond to OOM.
@@ Appendix F Long video memorisation task
Figure 9: Generalisation to longer sequences. Both models are trained using Imagenet pre-trained weights, on video sequences of $T=64$ frames to reconstruct the $16^{\text{th}}$ frame; during evaluation, the models receive sequences of $T=96$ frames.